OpenStudy (anonymous):

Find the slope of the tangent line to the curve \sqrt{1x +4y} + \sqrt{4xy} = 17 at the point ( 7,5 ). Specifically, I'm having trouble with the \sqrt{4xy} since it requires both product and chain rules (not sure which to do first and how).

hartnn (hartnn):

glad that you specified, whenever in doubt, try substitution, u= 4xy \((\sqrt{4xy})'=(\sqrt{u})'=1/2\sqrt{u}\times (4xy)'\) now product rule for (4xy)' don't forget to resubstitute back for 'u'

OpenStudy (anonymous):

sry, i don't understand, this is what I have so far: (1/2)(x+4y)^(-1/2) * (1+4(dy/dx)) + (1/2)[(4y)(4x(dy/dx))]^(-1/2)

hartnn (hartnn):

regarding 'which to do first ' do, chain rule first, so, there'll be 1/2 (4xy)^(-1/2) [....here product rule for (4xy)'....]

OpenStudy (anonymous):

(1/2)(x+4y)^(-1/2) * (1+4(dy/dx)) + (1/2)(xy)^(-1/2) * (4y)(4x(dy/dx)) Here is what I have after those steps, I'm still uncertain what to do next

hartnn (hartnn):

product rule for (4xy)' = 4 [y + x dy/dx] right ? so, 2nd term = 1/2 (4xy)^(-1/2) 4[y + x dy/dx] got this ? 1st term is just fine.

OpenStudy (anonymous):

more specifically, i mean im not sure what to do with the first terms, since i need to bring all the y primes on one side and the rest to the other (so i can plug in points (7,5). Im just not sure about how to deal with (1/2)(x+4y)^(-1/2) and the steps afterwards.

hartnn (hartnn):

you can plug in 7,5 before isolating y' also, this will simplify isolating y' afterwards.